OpenStudy (crashonce):

Trig @ganeshie8 @ikram002p

OpenStudy (crashonce):

\[\sin \frac{ \pi }{ 9}\sin \frac{ 2\pi }{ 9}\sin \frac{ 3\pi }{ 9}\sin \frac{ 4\pi }{ 9}= \frac{ 3 }{16 }\] Prove this

OpenStudy (crashonce):

@Nurali @ShadowLegendX

OpenStudy (ikram002p):

ok , which angle you know its sin ?

OpenStudy (crashonce):

Thats the question im not sure what u mean

OpenStudy (ikram002p):

well , this could be easily shown from calculator :o but if you wanna use identity , there is special angels would be its sin /cos given / known to evaluate this

OpenStudy (crashonce):

ty

OpenStudy (akashdeepdeb):

Did you prove it? Do you still need help?

OpenStudy (ikram002p):

:o ty for what did u got it ?